Prove that t75%/t50%=2 for a first order reactionSub chemistry model paper

Grade:12

2 Answers

Sandeep singh
6 years ago
T75%=2.303/k*log100/100-75. =2.303/k*log4=2.303/k*0.60205999. T50%=2.303/k*log100/100-50. =2.303/k*log2=2.303/k*0.30103. T75%/T50%=0.60205999/0.30103=1.999=2
Deepak kumar
4 years ago
 
T75%=2.303/k*log100/100-75. =2.303/k*log4=2.303/k*0.60205999.
T50%=2.303/k*log100/100-50. =2.303/k*log2=2.303/k*0.30103.
T75%/T50%=0.60205999/0.30103=1.999=2